1. The indicated function y1(x) is a solution of the given differential equation. Use reduction of order or formula (5) in Section 4.2,

y2 = y1(x) ∫ e^−∫P(x) dx / y upper 2 and lower 1 (x) dx

as instructed, to find a second solution y2(x).
y'' − 4y' + 4y = 0; y1 = e^2x

y2 = _____
1
Expert's answer
2020-06-30T16:19:19-0400

Solution:

y1(x)=e2x,

P(x)=-4,

"\int" x0P(x')dx'="\int"x0 (-4)dx'=-4x'|x0=-4x,

y2(x)=y1(x)"\int"e^(- "\int" x0 (-4)dx')/y12(x)dx=e2x"\int" e4x/e4xdx=e2x"\int" dx=e2x(x+C),

Let C=0, then y2(x)=xe2x

Answer: y2(x)=xe2x.